OpenStudy (clamin):

PLEASE HELP!! A ladder, 500 cm long, leans against a building. If the angle between the ground and the ladder is 57 degrees, how far from the wall is the bottom of the ladder? Round the answer to the nearest tenth.

OpenStudy (jadzia):

 Apply Trig Ratios: \(\sf \Large tan\ \theta=\frac{opposite}{adjacent}\) \(\sf \Large cos\ \theta=\frac{adjacent}{hypotenuse}\) \(\sf \Large sin\ \theta=\frac{opposite}{hypotenuse}\)

OpenStudy (anonymous):

From trig rules : \[\cos = \frac{ adjacent }{ hypotenus }\] So using this and from the fig : \[\cos(57) = \frac{ x }{ 500 }\] which simply means that x = 500 cos(57) = 272.3195175 which are able to round it with required Let me know if you got it :)